A force of 200N makes an angle θ and has a scalar y-component of 30N. Find both the scalar x-component of force and the angle θ. What displacement at 75
1
Expert's answer
2021-09-29T14:08:07-0400

"F_y=200\cdot \sin\theta\to \sin\theta=F_y/200=30/200=" 0.15


"\theta =8.6°"


"F_x=200\cdot \cos8.6=197.75\ (N)"